A pendulum oscillates 40 times is 4 seconds. Find its time period and frequency.

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

Number of oscillations = 40
Total time = 4 seconds
Time taken to complete one oscillation = 4/40 = 0. 1 second
So time period = 0.1 second
Frequent = 1/Time period = 1/0.1 = 10 Hz

Thanks